Circular No. 7446

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cs IN MCG +7-34-15 T. Puckett, Mountain Town, GA, reports the discovery of an appar-

ent supernova (mag 17.8) on an unfiltered CCD frame (limiting mag 20.5) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on June 30.25 UT. The candidate is located at α = 16^{h}23^{m}32^{s}.24, δ = +39^{o}07′29′′.1 (equinox 2000.0), which is 5′′.5 west and 11′′.9 north of the center of MCG +7-34-15. The new object was also present on an unfiltered CCD frame taken by Puckett on July 1.13. The candidate was also confirmed by J. Newton, Newton Observatory B & B, Osoyoss, BC, on unfiltered CCD images taken with a 0.40-m reflector on July 1.23. The candidate does not appear on Palomar Sky Survey images taken on 1993 Apr. 30, 1989 May 31 and 1953 May 19 (limiting mag ∼ 21.0, 21.0 and 20.0, respectively).

4U 2206+54 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association; R. Remillard, Massachusetts Institute of Technology (MIT); and A. Peele, University of Melbourne, report on behalf of the RXTE ASM team at GSFC and MIT: “We have found that the RXTE ASM lightcurve of the Be star and x-ray source 4U 2206+54 had previously been made using a position differing by 0^{o}.5 from the correct location of α = 22^{h}07^{m}57^{s}.0, δ = +54^{o}31′06′′ (equinox 2000.0; Steiner et al. 1984, Ap.J. 280, 688). We have now produced and analyzed a revised lightcurve. The mean source flux is ∼ 5 mCrab, and we find quasisinusoidal modulation at a period of 9.570 ± 0.004 days with a semiamplitude of ∼ 1.5 mCrab. There was a maximum on 1998 July 12.1 ± 0.2 UT. If this periodicity is due to orbital motion, then it would be one of the shortest known for a Be-star x-ray source. However, the relatively low luminosity (< 10^{35} ergs/s at 2.5 kpc) would run counter to the general trend for short-orbital-period Be/neutron-star systems to have higher luminosities. Observations at other wavelengths, including optical spectroscopy and photometry, are encouraged.”

SUNGRAZING COMETS Further to IAUC 7445, it should be noted that comets C/1998 J4

and C/2000 L4 showed tails, while C/1998 J3 did not. C/2000 L5 was independently reported (mag 4–5) to the Central Bureau by A. Mimeev. The correct transliteration of the name of the observer of C/2000 M1 is P. Shkreby.

2000 July 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden